### v1

@@ 7 Conclusions
This paper presents our solution for long-duration inspection: a tethered marsupial system connecting a UAV and a UGV. In it, we detail the hardware and software configurations that allow autonomous navigation in UAV inspection tasks. Notably, the system achieved over one hour of flight time—and up to a maximum of two hours—thanks to the power supplied by the batteries onboard the UGV. Therefore, the proposed marsupial system successfully extends the operational capabilities of the UAV.
The experimental results obtained across three different scenarios validate the robustness and operational capabilities of our system in real-world conditions, delivering reliable inspection performance in challenging, GNSS-denied environments. Specifically, a series of tests conducted at the University Pablo de Olavide confirmed the system’s ability to carry out autonomous, long-duration inspection missions. Notably, our implementation of the DLL localization system provided real-time pose estimations with centimeter-level accuracy, outperforming state-of-the-art methods. Furthermore, our path tracking system successfully followed the reference trajectories, maintaining maximum tracking errors below ten centimeters. As a result, the system was able to detect simulated defects with centimeter-level accuracy during the emulated inspection test.
@@
Future work will focus on refining the power system to further extend the operational range. Additionally, we are developing a continuous control system within the TCM module to enable smooth coordination between platforms by adjusting the tracking velocity of each system, moving away from the current Stop & Go approach. Moreover, we are interested in expanding the inspection capabilities of the system by integrating additional sensors required for in-depth inspection tasks, such as radar and IR cameras. However, the UAV’s payload limitations may prevent it from carrying all sensors simultaneously. To address this, we are taking two key actions. First, we are designing modular sensor attachments that allow easy reconfiguration of the UAV based on the specific requirements of each inspection mission. Second, Elistair offers a tethered station which includes an optical fiber communication link. In this way, we could remove the PC onboard the UAV, making room for installing additional sensors.
Declarations
[…]
Funding This work was supported by the grants INSERTION PID2021-127648OB-C31 and RATEC PDC2022-133643-C21, funded by MCIN/AEI/ 10.13039/501100011033 and ”European Union NextGenerationEU/PRTR”.
[…]
Conflict of interest/Competing interests The authors declare that they have no conflicts of interest or competing interests.
[…]
Ethics Approval Not Applicable.
[…]
Consent for publication Not Applicable.
[…]
Consent to participate Not Applicable.
[…]
Data availability
Data will be available upon reasonable request. Also, the code is available on the GitHub organization "robotics-upo" [32].
[…]
Authors’ Contributions Simón Martínez-Rozas and David Alejo: Conceptualization, data curation, software, investigation, methodology, analysis, experiments, writing original draft. José Javier Carpio: Mechanical development and maintenance of the robot, experiments, writing-review and editing. Fernando Caballero and Luis Merino: Conceptualization, validation, resources, supervision, funding acquisition, project management, writing-review and editing.
[…]
@@ Acknowledgements
We would like to thank the staff of the Service Robotics Laboratory at the Universidad Pablo de Olavide for their unselfish help during the experiments. We would also like to give special thanks to Sara Victoria Hernández Díaz for suggesting different experimental scenarios for structural inspection and granting us access to them.

### v2

@@ 7 Conclusions
This paper presents our solution for long-duration inspection: a tethered marsupial system connecting a UAV to a UGV. In it, we detail the hardware and software configurations that allow autonomous navigation in UAV inspection tasks, as demonstrated in several field experiments. Notably, we have demonstrated that the proposed system has a flight duration of over two hours thanks to the power supplied by the batteries onboard the UGV. Therefore, the proposed marsupial system successfully extends the operational capabilities of the UAV.
The proposed tethered marsupial system can successfully perform autonomous inspection tasks, detecting targets such as markers without being hindered by the tether. The inspection tasks are carried out with high accuracy both in terms of the localization of the marsupial robot and in the detection of the markers.
Future Work
Future work will focus on refining the power system to further extend the operational range. Additionally, we are developing a continuous control system within the TCM to enable smooth coordination between platforms by adjusting the tracking velocity of each system, moving away from the current stop-and-go synchronization.
@@
Moreover, we aim to expand the inspection capabilities of the system by integrating additional sensors required for advanced inspection tasks, such as radar and IR cameras. However, the UAV’s payload limitations may prevent it from carrying all sensors simultaneously. To address this, we are pursuing two complementary approaches. First, we are designing modular sensor attachments that allow easy reconfiguration of the UAV based on the specific requirements of each inspection mission. Second, Elistair offers a tethered station that includes an optical fiber communication link. In this way, we could remove the computer onboard the UAV, making room for additional sensors.
Another area of interest is the investigation of the applicability of our system in outdoor scenarios, or even in mixed scenarios that encompass both indoor and outdoor areas. These mixed scenarios pose significant challenges for the localization system, as they involve partial GNSS coverage and require ensuring a smooth transition between GNSS and GNSS-denied localization modes. In addition, the environmental factors such as wind can adversely affect the performance of the control system.
Finally, we aim to also use this approach in emergency response situations. To this end, we plan to modify the traction type of our UGV platform to enable safe and reliable navigation over uneven terrain. Additionally, we plan to include a multi-robot SLAM method to perform autonomous exploration of unknown scenarios.
All the data captured in the experiments is available at https://robotics.upo.es/datasets/marsupial, accessed on 3 Nov 2025.
@@ Acknowledgements
We would like to thank the staff of the Service Robotics Laboratory at UPO for their unselfish help during the experiments. We would also like to give special thanks to Sara Victoria Hernández Díaz for suggesting different experimental scenarios for structural inspection and granting us access to them.
This work was partially supported by the grants INSERTION PID2021-127648OB-C31 and RATEC PDC2022-133643-C21, funded by MCIN/AEI/ 10.13039/501100011033 and “European Union NextGenerationEU/PRTR”.